OpenStudy (anonymous):

eliminate the parameter and write the equation of the corresponding rectangular equation. x= t-1 y= t/t-1

OpenStudy (anonymous):

\(x(t)=t-1,\;y(t)=\frac{t}{t-1}\) Since x = t - 1, you have t = x + 1. By substitution, you get \[y=\frac{x+1}{x}=1+\frac{1}{x}\]